Can someone help me to calculate Cash Flow statement?

A. Income Statement Data for 20XX:

Units produced and sold = 420

Sales ($80 per unit selling price) = $33600

Cost of goods sold ($30 per unit, all variable costs) = $12600

Labor = $0 (Mr. and Mr. Lee were the only ones working and did not pay themselves)

Advertising fees =$2000

Bank fees = $150

Phone/internet = $1200

Shipping ($3 per unit) = $1260

Utilities = $900

Office supplies = $800

Interest expense on note payable = $350

Depreciation expense (straight line) = $800

My answer:

Particulars Amount
Sales $      33,600.00
Cost of Goods Sold $      12,600.00
Gross Profit $      21,000.00
Less Operating and Admin. Expenses
Advertising $        2,000.00
Bank Fees $            150.00
Phone/Internet $        1,200.00
Shipping $        1,260.00
Utilities $            900.00
Office Suppies $            800.00
EBITDA $      14,690.00
Depriciation $            800.00
EBIT $      13,890.00
Less
Interest on Notes Payable $            350.00
EBT $      13,540.00
Less Tax at 26% $       (3,520.40)
Net Profit $      10,019.60

Now I need to calculate cash flow

B. Proceeds from sale of equipment = $3000. The equipment originally cost $1000 and had accumulated depreciation of $200.

Purchase of equipment = $1600 (The machine is purchased on the last day of 20XX so no depreciation expense is recorded.)

Repayment of note payable = $5000

Consider any data relevant from the income statement.
Answer #1

Cashflow statement

Cashflow from operating activity = EBIT ( 1- Tax) + Depreciation - Capex +/- change in net working capital

= 13890 * ( 1 - 0.26) + 200 - 1600 - 5000 = 3878.6

Cashflow from Investing activity = 3000 (sale of equipment)

Cashflow = 3878.6 + 3000 = 6878.6